# Locally homogeneous Axiom A flows II: geometric structures for Anosov subgroups

### v2

@@ 3.7.1. Jordan projections in $G$
Let us fix a restricted root space decomposition as in Section 3.4, in particular we fix an Iwasawa decomposition $G=KAN$ which, in contrast to the Jordan decomposition, involves making choices. Then a group element being elliptic (resp. hyperbolic, unipotent) means that it is conjugate to an element in $K$ (resp. $A$, $N$).
@@ 3.7.3. Class functions on $G$
The Jacobian class functions of a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ are related by
@@ 3.8. Linear actions on tangent spaces of flag manifolds
Every element $X\in\mathfrak{g}$ induces fundamental vector fields on ${\mathcal{F}}^{\pm}$ and ${{\mathcal{F}}^{\pitchfork}}$, which we shall all denote by $\overline{X}$, via the formula $\overline{X}(x)=\left.\frac{d\,}{dt}\right|_{t=0}\exp(tX)\cdot x$. The kernel of the evaluation map
@@
Note that these isomorphisms swap the $+$ and -- signs in the upper index, which will be important to keep track of when we apply them.
@@
[…]
@@
Let ${\mathcal{F}}$ be a flag manifold, $\mathbf{x}\in{\mathcal{F}}$, and $g\in G$. We say that $\mathbf{x}$ is an attracting fixed point of $g$ if $g\in\mathrm{Stab}_{G}(\mathbf{x})$ and the linear action of $g$ on $T_{\mathbf{x}}{\mathcal{F}}$ given by the derivative of the $G$-action has spectral radius $<1$.
@@
Definition 3.38.
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ and the transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$. An element $g\in G$ is ${{\mathcal{F}}^{\pitchfork}}$-proximal if $g$ is ${\mathcal{F}}^{+}$-proximal (which by (3.34) is equivalent to $g^{-1}$ being ${\mathcal{F}}^{-}$-proximal) and the attracting fixed points $g^{+}\in{\mathcal{F}}^{+}$, $g^{-}\in{\mathcal{F}}^{-}$ of $g$ and $g^{-1}$ satisfy $(g^{+},g^{-})\in{{\mathcal{F}}^{\pitchfork}}$. In this case, $g^{-}$ is called the repelling fixed point of $g$.
@@
Let ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ be a pair of opposite flag manifolds corresponding to subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively. Denote the transverse flag spaces by ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ and ${\mathcal{F}}_{\alpha}^{\pitchfork}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$ for $\alpha\in\Theta$. Then any ${{\mathcal{F}}^{\pitchfork}}$-proximal element $g\in G$ with attracting/repelling fixed points $(\mathbf{x}^{+},\mathbf{x}^{-})\in{{\mathcal{F}}^{\pitchfork}}$ is also ${\mathcal{F}}^{\pitchfork}_{\alpha}$-proximal with attracting/repelling fixed points $(\mathbf{x}^{+}_{\alpha},\mathbf{x}^{-}_{\alpha})\in{\mathcal{F}}^{\pitchfork}_{\alpha}$ for all $\alpha\in\Theta$, where $\mathbf{x}^{+}_{\alpha}=\Pi^{{\mathcal{F}}^{+}}_{\alpha}(\mathbf{x}^{+})$ and $\mathbf{x}^{-}_{\alpha}=\Pi^{{\mathcal{F}}^{-}}_{\iota_{o}(\alpha)}(\mathbf{x}^{-})$ are the images under the unique projections introduced in (3.15).
@@
Let $\mathcal{F}_{\alpha}$ be a simple flag manifold and ${\mathcal{F}}^{\pitchfork}_{\alpha}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$ the corresponding transverse flag space. For any $\mathcal{F}^{\pitchfork}_{\alpha}$-proximal element $g\in G$, we have
@@
Just as in the proof of Lemma 3.39, $\mathcal{F}^{\pitchfork}_{\alpha}$-proximality means that we can work in the context of a fixed restricted root space decomposition and assume that $g=\exp(X)$ for some $X\in{\bf\mathfrak{a}}^{+}$, in which case $\mathcal{P}_{\alpha}(g)=\alpha(X)$ and the derivative $d_{\mathbf{x}^{-}}g$ is conjugate to the action of $\mathrm{Ad}(\exp(X))=\exp(\operatorname{ad}_{X})$ on the nilpotent radical $\mathfrak{n}=\bigoplus_{\beta\in\Sigma_{\{\alpha\}}^{+}}{\bf\mathfrak{g}}_{\beta}$, so we find
@@ 4.3. Homogeneous affine line bundles and additive characters
There is a bijective correspondence between Lie group homomorphisms $b:H\to\mathbb{R}$ (considered up to precomposition by an isomorphism) and (equivalence classes of) $G$-equivariant affine line bundles over $G/H$.∎
@@ 4.4. The period function of a $G$-equivariant affine line bundle
The equivalence between (1) and (3) is Lemma 4.2, and it is straightforward to see that (1$\Rightarrow$(2$\Rightarrow$(3). ∎
@@
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ and the transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$. The Jacobian characters of Levi subgroups can be combined into the period function
@@
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ and their corresponding subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$, and let ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$. The functions $(\mathbf{J}_{{{\mathcal{F}}^{\pitchfork}}}^{\alpha})_{\alpha\in\Theta}$ form a basis of $\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$.
@@ 5.1. General definitions and properties
Let $E$ be a finite dimensional real vector space of dimension $d>0.$ Let $\mathcal{B}_{E}$ be the set of ordered bases of $E$, which is the unique open $\mathrm{GL}(E)$-orbit in the product of $d$ copies of $E$
@@
A non-degenerate symmetric bilinear form $\varphi:E\times E\to\mathbb{R}$ produces a positive $s$-density $\mathrm{vol}_{\varphi}^{s}\in\mathscr{D}^{s}_{>0}(E)$ defined by $\mathrm{vol}_{\varphi}^{s}(\mathbf{e})=|\mathrm{det}\,[\varphi]_{\mathbf{e}}|^{s/2}$, where the $(d\times d)$-Gram matrix $[\varphi]_{\mathbf{e}}$ has coefficients $\varphi(e_{i},e_{j})$ for $\mathbf{e}=(e_{1},\dots,e_{d})$.
@@ 5.2.1. Density bundles
of positive $s$-densities on ${\mathbb{F}}$. When ${\mathbb{F}}$ carries a $G$-invariant pseudo-Riemannian metric (such is the case of a transverse flag space ${\mathbb{F}}={\mathcal{F}}^{\pitchfork}$, as seen in Section 3.3), the $G$-equivariant affine line bundle $\mathscr{D}^{s}_{>0}{{\mathbb{F}}}$ has a $G$-equivariant section, so it is $G$-equivariantly trivializable.
When ${\mathbb{F}}={\mathcal{F}}$ is a non-trivial flag manifold and $s\neq 0$, then $\mathscr{D}^{s}_{>0}{{\mathcal{F}}}$ is a homogeneous affine line bundle. Given a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$, the period functions of the density bundles $\mathscr{D}^{s}_{>0}{\mathcal{F}}^{\pm}$ are related to Jacobian characters of Levi subgroups:
@@
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$, suppose that $g\in G$ has a pair of transverse fixed points $(\mathbf{x}^{+},\mathbf{x}^{-})\in{\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ and write $L=\mathrm{Stab}_{G}(\mathbf{x})$ for the corresponding Levi subgroup. Then one has
@@
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$, let $\sigma$ be a section of $\mathscr{D}^{s}_{>0}{\mathcal{F}}^{-}$ for some $s\in\mathbb{R}$, $g\in G$ an ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$-proximal element with attracting fixed point ${\mathbf{x}}^{+}\in{\mathcal{F}}^{+}$ and fix a basis $(X_{1},\dots,X_{n})$ of the nilpotent radical $\mathfrak{n}$ of the Lie algebra of its stabilizer. Then for any $\mathbf{x}\in{\mathcal{F}}^{-}$ transverse to ${\mathbf{x}}^{+}$, one has
@@ 5.2.2. Multidensity bundles
[…]
@@
By Lemma 3.40, each pair $(\mathbf{x}^{+}_{\alpha},\mathbf{x}^{-}_{\alpha})$ is transverse, so the claim follows from Corollary 5.7. Note that the change of signs present in Corollary 5.7, which results from Remark 3.1, is reflected here in the appearance of $g^{-1}$ in the $+$ case and $g$ in the -- case. ∎
@@ 6. Flow spaces
With the preparations from Sections 4 and 5 at hand, it is now time to introduce the desired flow spaces. The idea is to create a multidensity analog of the hyperbola branch $\{(x^{+},x^{-})\in\mathbb{R}_{>0}^{2}\,|\,x^{+}\cdot x^{-}=1\}$. To this end, we need to pullback both $\mathbf{s}$-multidensity bundles $\mathscr{D}^{\mathbf{s}}{\mathcal{F}}^{+}_{\bullet}$ and $\mathscr{D}^{\mathbf{s}}{\mathcal{F}}^{-}_{\bullet}$ to the transverse flag space ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ where a pairing can be defined thanks to the pseudo-Riemannian structure.
@@ 6.1. Pairings of multidensities on a transverse flag space
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ with corresponding subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively, and a weight vector $\mathbf{s}\in\mathbb{R}^{\Theta}$. By using the projections
${{\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}}$${{\mathcal{F}}^{+}}$${{\mathcal{F}}^{-}}$$\scriptstyle{p_{{\mathcal{F}}^{+}}}$$\scriptstyle{p_{{\mathcal{F}}^{-}}}$
@@ 6.2. The multidensity flow
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ with corresponding subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively, and a weight vector $\mathbf{s}\in\mathbb{R}^{\Theta}$.
The flow space $\mathbb{L}^{\mathbf{s}}_{\Theta}$ of ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ with parameter $\mathbf{s}$ is the $G$-equivariant affine line bundle over ${{\mathcal{F}}^{\pitchfork}}$ given by
@@
$[{\mathbb{L}}^{\mathbf{s}}_{\Theta}]=\big{[}p_{{\mathcal{F}}^{-}}^{\ast}\big{(}\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}\big{)}\big{]}=\big{[}p_{{\mathcal{F}}^{+}}^{\ast}\big{(}\overline{\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{+}_{\bullet}}\big{)}\big{]}.$ (6.4)
@@ 7. Anosov subgroups
The proximality class functions $\mathcal{P}_{\alpha}:G\to\mathbb{R}_{\geq 0}$ introduced in Definition 3.25 for the simple flag manifolds ${\mathcal{F}}_{\alpha}$, $\alpha\in\Delta$ provide an intrinsic way of stating Kassel-Potrie’s definition of Anosov subgroups from [KP22], in which we can flexibly switch between the different interpretations of the set $\Delta$ explained in Section 3.6.
@@
Given a subset $\Theta\subset\Delta$, a word-hyperbolic discrete subgroup $\Gamma<G$ is called $\Theta$-Anosov if there are constants $C,c>0$ such that
$\mathcal{P}_{\alpha}(\gamma)\geq c|\gamma|_{\infty}-C$
holds for all $\alpha\in\Theta$ and $\gamma\in\Gamma$.
Here $|\cdot|_{\infty}$ is the stable length with respect to some finite generating set of $\Gamma$, the choice of which does not matter in the definition (see Section 2.2 for more details and references).
Let ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ be a pair of opposite flag manifolds associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ their transverse flag space, and let $\Gamma<G$ be a $\Theta$-Anosov subgroup. We denote by $\partial_{\infty}\Gamma$ its Gromov-boundary and write
$\partial_{\infty}\Gamma^{(2)}:=\{(x,y)\in\partial_{\infty}\Gamma\times\partial_{\infty}\Gamma\,|\,x\neq y\}.$
The following consequences of the $\Theta$-Anosov property of $\Gamma$ are proved in [KP22, GGKW17, KLP17, GW12]:
@@
    The boundary maps preserve the convergence properties of the $\Gamma$-action on $\partial_{\infty}\Gamma$: For any unbounded sequence $\gamma_{n}\in\Gamma$ with boundary limit points $\gamma_{+}=\lim_{n\to+\infty}\gamma_{n}\in\partial_{\infty}\Gamma$ and $\gamma_{-}=\lim_{n\to+\infty}\gamma_{n}^{-1}\in\partial_{\infty}\Gamma$, there is a subsequence $\gamma_{n_{k}}$ for which the $\Gamma$-actions on ${\mathcal{F}}^{+}$ and ${\mathcal{F}}^{-}$ obey the following dynamics as $k\to+\infty$:
@@
      $\gamma_{n_{k}}\cdot\mathbf{x}^{+}\rightarrow\xi^{+}(\gamma_{+})$ for all $\mathbf{x}^{+}\in{\mathcal{F}}^{+}$ with $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$;
@@
      $\gamma_{n_{k}}^{-1}\cdot\mathbf{x}^{+}\rightarrow\xi^{+}(\gamma_{-})$ for all $\mathbf{x}^{+}\in{\mathcal{F}}^{+}$ with $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{+})$;
@@
      $\gamma_{n_{k}}\cdot\mathbf{x}^{-}\rightarrow\xi^{-}(\gamma_{+})$ for all $\mathbf{x}^{-}\in{\mathcal{F}}^{-}$ with $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$;
@@
      $\gamma_{n_{k}}^{-1}\cdot\mathbf{x}^{-}\rightarrow\xi^{-}(\gamma_{-})$ for all $\mathbf{x}^{-}\in{\mathcal{F}}^{-}$ with $\xi^{+}(\gamma_{+})\pitchfork\mathbf{x}^{-}$;
@@
We combine the boundary maps $\xi^{+},\xi^{-}$ to the map
@@
The images of $\xi^{\pm}$ and $\xi^{\pitchfork}$ are the limit sets in ${\mathcal{F}}^{\pm}$ and ${{\mathcal{F}}^{\pitchfork}}$, respectively, denoted by
@@
[…]

### v3

@@ 3.7.1. Jordan projections in $G$
Let us fix a restricted root space decomposition as in Section 3.4, in particular we fix an Iwasawa decomposition $G=KAN$ which, in contrast to the Jordan decomposition, involves making choices. Then a group element being elliptic (resp. hyperbolic, unipotent) means that it is conjugate to an element in $K$ (resp. $A$, NN).
@@ 3.7.3. Class functions on $G$
The Jacobian class functions of a pair of opposite flag manifolds $\mathcal{F}$$^{+}$$\mathcal{F}$$^{-}$ are related by
@@ 3.8. Linear actions on tangent spaces of flag manifolds
[…]
@@
Note that these isomorphisms swap the ++ and - signs in the upper index, which will be important to keep track of when we apply them.
@@
Let $\mathcal{F}$$^{+}$$\mathcal{F}$$^{-}$ be a pair of opposite flag manifolds. Consider $\mathbf{x}=(\mathbf{x}^{+},\mathbf{x}^{-})\in{\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$, let $L=\mathrm{Stab}_{G}(\mathbf{x})$ denote its stabilizer and $\mathfrak{n}$ the nilpotent radical of the Lie algebra of the stabilizer of $\mathbf{x}^{+}\in{\mathcal{F}}^{+}$. Then for any $g\in L$, we have
@@
Let $\mathcal{F}$} be a flag manifold, $\mathbf{x}\in{\mathcal{F}}$, and $g\in G$. We say that $\mathbf{x}$ is an attracting fixed point of $g$ if $g\in\mathrm{Stab}_{G}(\mathbf{x})$ and the linear action of $g$ on $T_{\mathbf{x}}{\mathcal{F}}$ given by the derivative of the $G$-action has spectral radius <1<1.
@@
Remark 3.2.
Consider a pair of opposite flag manifolds $\mathcal{F}$$^{+}$$\mathcal{F}$$^{-}$, and an $\mathcal{F}^{+}$-proximal element $g\in G$. Denoting by $g^{+}\in\mathcal{F}^{+}$ its attracting fixed point and by $g^{-}\in\mathcal{F}^{-}$ its repelling fixed point, i.e. the attracting fixed point of $g^{-1}$, we have that $g^{+}\pitchfork g^{-}$.
@@
[…]
@@
Let $\mathcal{F}_{\alpha}$ be a simple flag manifold and ${\mathcal{F}}^{\pitchfork}_{\alpha}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$ the corresponding transverse flag space. For any $\mathcal{F}_{\alpha}$-proximal element $g\in G$, we have
@@
Just as in the proof of Lemma 3.38, $\mathcal{F}_{\alpha}$-proximality means that we can work in the context of a fixed restricted root space decomposition and assume that $g=\exp(X)$ for some $X\in{\bf\mathfrak{a}}^{+}$, in which case $\mathcal{P}_{\alpha}(g)=\alpha(X)$ and the derivative $d_{\mathbf{x}^{-}}g$ is conjugate to the action of $\mathrm{Ad}(\exp(X))=\exp(\operatorname{ad}_{X})$ on the nilpotent radical $\mathfrak{n}=\bigoplus_{\beta\in\Sigma_{\{\alpha\}}^{+}}{\bf\mathfrak{g}}_{\beta}$, so we find
@@ 4.3. Homogeneous affine line bundles and additive characters
There is a bijective correspondence between Lie group homomorphisms $b:H\to\mathbb{R}$ (considered up to precomposition by an isomorphism) and (equivalence classes of) $G$-equivariant affine line bundles over G/HG/H.∎
@@ 4.4. The period function of a $G$-equivariant affine line bundle
[…]
@@
[…]
@@
[…]
@@ 5.1. General definitions and properties
Let $E$ be a finite dimensional real vector space of dimension $d>0.$ Let $\mathcal{B}_{E}$ be the set of ordered bases of $E$, which is the unique open $\mathrm{GL}(E)$-orbit in the product of $d$ copies of $E$ .
@@
A non-degenerate symmetric bilinear form $\varphi:E\times E\to\mathbb{R}$ produces a positive $s$-density $\mathrm{vol}_{\varphi}^{s}\in\mathscr{D}^{s}_{>0}(E)$ defined by $\mathrm{vol}_{\varphi}^{s}(\mathbf{e})=|\mathrm{det}\,[\varphi]_{\mathbf{e}}|^{s/2}$, where the $(d\times d)$-Gram matrix $\varphi]_{\mathbf{e}}$ has coefficients $\varphi(e_{i},e_{j})$ for $\mathbf{e}=(e_{1},\dots,e_{d})$.
@@ 5.2.1. Density bundles
[…]
When $\mathbb{F}$$\mathcal{F}$} is a non-trivial flag manifold and $s\neq 0$, then $\mathscr{D}^{s}_{>0}{{\mathcal{F}}}$ is a homogeneous affine line bundle. Given a pair of opposite flag manifolds $\mathcal{F}$$^{+}$$\mathcal{F}$$^{-}$, the period functions of the density bundles $\mathscr{D}^{s}_{>0}{\mathcal{F}}^{\pm}$ are related to Jacobian characters of Levi subgroups:
@@
[…]
@@
[…]
@@ 5.2.2. Multidensity bundles
$T{\mathcal{F}}_{\alpha_{1}}$$\cdots$$T{\mathcal{F}}_{\alpha_{k}}$${\mathcal{F}}_{\alpha_{1}}$$\cdots$${\mathcal{F}}_{\alpha_{k}}$$\mathcal{F}$$\scriptstyle{\Pi_{\alpha_{1}}^{\mathcal{F}}}$$\scriptstyle{\Pi^{\mathcal{F}}_{\alpha_{k}}}$
@@
By Lemma 3.39, each pair $(\mathbf{x}^{+}_{\alpha},\mathbf{x}^{-}_{\alpha})$ is transverse, so the claim follows from Corollary 5.7. Note that the change of signs present in Corollary 5.7, which results from Remark 3.1, is reflected here in the appearance of $g^{-1}$ in the ++ case and $g$ in the - case. ∎
@@ 6. Flow spaces
[…]
@@ 6.1. Pairings of multidensities on a transverse flag space
Consider a pair of opposite flag manifolds $\mathcal{F}$$^{+}$$\mathcal{F}$$^{-}$ with corresponding subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively, and a weight vector $\mathbf{s}\in\mathbb{R}^{\Theta}$. By using the projections
[…]
@@ 6.2. The multidensity flow
Consider a pair of opposite flag manifolds $\mathcal{F}$$^{+}$$\mathcal{F}$$^{-}$ with corresponding subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively, and a weight vector $\mathbf{s}\in\mathbb{R}^{\Theta}$.
[…]
@@
${\mathbb{L}}^{\mathbf{s}}_{\Theta}]=\big[p_{{\mathcal{F}}^{-}}^{\ast}\big(\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}\big)\big]=\big[p_{{\mathcal{F}}^{+}}^{\ast}\big(\overline{\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{+}_{\bullet}}\big)\big].$ (6.4)
@@ 7. Anosov subgroups
[…]
@@
A hyperbolic discrete subgroup $\Gamma<G$ is called $\Theta$-Anosov if
@@
    The boundary maps preserve the convergence properties of the $\Gamma$-action on $\partial_{\infty}\Gamma$: For any unbounded sequence $\gamma_{k}\in\Gamma$ with boundary limit points $\gamma_{+}=\lim_{k\to+\infty}\gamma_{k}\in\partial_{\infty}\Gamma$ and $\gamma_{-}=\lim_{k\to+\infty}\gamma_{k}^{-1}\in\partial_{\infty}\Gamma$, the $\Gamma$-actions on $\mathcal{F}$$^{+}$ and $\mathcal{F}$$^{-}$ obey the following dynamics as $k\to+\infty$:
@@
      $\gamma_{k}\cdot\mathbf{x}^{+}\rightarrow\xi^{+}(\gamma_{+})$ for all $\mathbf{x}^{+}\in{\mathcal{F}}^{+}$ with $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$;
@@
      $\gamma_{k}^{-1}\cdot\mathbf{x}^{+}\rightarrow\xi^{+}(\gamma_{-})$ for all $\mathbf{x}^{+}\in{\mathcal{F}}^{+}$ with $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{+})$;
@@
      $\gamma_{k}\cdot\mathbf{x}^{-}\rightarrow\xi^{-}(\gamma_{+})$ for all $\mathbf{x}^{-}\in{\mathcal{F}}^{-}$ with $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$;
@@
      $\gamma_{k}^{-1}\cdot\mathbf{x}^{-}\rightarrow\xi^{-}(\gamma_{-})$ for all $\mathbf{x}^{-}\in{\mathcal{F}}^{-}$ with $\xi^{+}(\gamma_{+})\pitchfork\mathbf{x}^{-}$;
@@
Remark 7.1.
This is not the original definition of $\Theta$-Anosov subgroups as given in [Lab06] or [GW12], but an equivalent characterisation proved in [KLP17, Theorem 1.1], where it is called asymptotically embedded. Limit maps satisfying (A1), (A2) and (A3) are unique.
Lemma 7.2.
Let $\Gamma<G$ be $\Theta$-Anosov, and consider a sequence $(\gamma_{k})$ in $\Gamma$ possessing boundary limits $\gamma_{+}=\lim_{k\to\infty}\gamma_{k},\gamma_{-}=\lim_{k\to\infty}\gamma_{k}^{-1}\in\partial_{\infty}\Gamma$. If $\gamma_{+}\neq\gamma_{-}$, then
$\lim_{k\to\infty}\mathcal{P}_{\alpha}(\gamma_{k})=+\infty~,\quad\forall\alpha\in\Theta\,.$
Proof.
By Lemma 2.5, we have that $|\gamma_{k}|_{\infty}\to+\infty$, and [GGKW17, Theorem 1.7] guarantees that $\mathcal{P}_{\alpha}(\gamma)\to+\infty$ as $|\gamma|_{\infty}\to+\infty$. ∎
For a hyperbolic group $\Gamma$, we write
$\partial_{\infty}\Gamma^{(2)}:=\{(z^{+},z^{-})\in\partial_{\infty}\Gamma\times\partial_{\infty}\Gamma\,|\,z^{+}\neq z^{-}\}.$
Let $\Gamma<G$ be a $\Theta$-Anosov subgroup, and combine the boundary maps $\xi^{+},\xi^{-}$ to the map
@@
[…]
@@
$\Gamma\text{ is }\Theta\text{-Anosov}\iff\Gamma\text{ is }\iota_{o}(\Theta)\text{-Anosov}.$ (7.2)